AGS Polarimeter Status

26 Feb 2008: Getting ready to put in new targets.

01 Nov 2007: New drive mechanism installed that should prevent or reduce the possibility of colliding targets.  New targets installed.  Cables for photodiode testing installed.  Waiting for new detector and new cables.

28 Sept 2007:  There are plans to pull ~50 cables from C-15 to the HITL room in order to test a new type of silicon detector for the AGS Polarimeter.  This work is scheduled over then next three weeks.


Scheduling Physicist: Mike Sivertz Last updated:
Email: sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528
Home Phone: 631-878-9020